# Physics-Informed Multi-Agent Reinforcement Learning for Distributed Multi-Robot Problems

### v1

@@ Abstract
The networked nature of multi-robot systems presents challenges in the context of multi-agent reinforcement learning. Centralized control policies do not scale with increasing numbers of robots, whereas independent control policies do not exploit the information provided by other robots, exhibiting poor performance in cooperative-competitive tasks. In this work we propose a physics-informed reinforcement learning approach able to learn distributed multi-robot control policies that are both scalable and make use of all the available information to each robot. Our approach has three key characteristics. First, it imposes a port-Hamiltonian structure on the policy representation, respecting energy conservation properties of physical robot systems and the networked nature of robot team interactions. Second, it uses self-attention to ensure a sparse policy representation able to handle time-varying information at each robot from the interaction graph. Third, we present a soft actor-critic reinforcement learning algorithm parameterized by our self-attention port-Hamiltonian control policy, which accounts for the correlation among robots during training while overcoming the need of value function factorization. Extensive simulations in different multi-robot scenarios demonstrate the success of the proposed approach, surpassing previous multi-robot reinforcement learning solutions in scalability, while achieving similar or superior performance (with averaged cumulative reward up to ×$\times 2$ greater than the state-of-the-art with robot teams ×$\times 6$ larger than the number of robots at training time).
@@ I Introduction
navigation sampling
Refer to caption Refer to caption
transport grassland
Refer to caption Refer to caption
Figure 1: Examples of scenarios addressed by our physics-informed multi-agent reinforcement learning approach. The scenarios cover a wide variety of cooperative/competitive behaviors and levels of coordination complexity.
@@
Extensive simulations in different multi-robot scenarios (Sec. VI) demonstrate that the combination of multi-agent reinforcement learning techniques and a physics-informed description of the system achieves scalable control policies with similar or superior performance as the state-of-the-art (Sec. VII).
This paper is an extension of a conference paper in [17]. This paper addresses multi-robot problems where only a reward specification of the task is given, whereas [17] considers that data from an expert controller is available. In that sense, [17] is more closely related to inverse reinforcement learning, while this paper considers a reinforcement learning setting. The proposed approach is also different. Specifically, we propose a novel multi-agent reinforcement learning approach where the control policy is conditioned on the interaction graph of the multi-robot system by exploiting a port-Hamiltonian formulation and a self-attention-based parameterization, while [17] reduces to a supervised learning approach. The neural networks that model the control policies are different as well, since the multi-agent reinforcement learning scenarios are continuous-time while the datasets in [17] are discrete-time. Finally, all experiments in this paper are new.
@@ II-B Multi-agent reinforcement learning for robotic problems
To solve the non-stationarity issue, the common approach is to factorize the value and/or policy functions. The most extended factorization, the so-called centralized-training decentralized-execution, departs from centralized-training reinforcement learning approaches [54, 55, 56] and considers that the agents act independently. To do so, the agents can learn an approximation of the other agents’ control policies [57]. The resulting control policies only use the individual observations; by neglecting communications, the performance barely scales with the number of robots. Another option is to provide each agent with a global estimator that predicts the trajectories of the whole multi-robot team [58] or to implicitly learn to coordinate with other agents through the value function [53]. In mean field approaches [59], the multi-agent game is reduced to an interaction between an agent and the average of the other agents, which is not practical. To overcome these issues, we propose a new formulation of the policy function that is distributed by design and captures the underlying interaction graph that describes the multi-robot system.
Other works focus on more complex factorizations that account for the local interactions among neighboring agents in the value function. For instance, the value function can be approximated as dependent on the neighbors only [60]. Either from analytical [33, 14] or learned [61, 62, 63] factorizations, the learned control policies are still restricted to the same team size used during training because the policies do not explicitly consider communication. By using a port-Hamiltonian description of the system, our approach avoids to approximate the reward, value or policy functions by any factorization, thus fully exploiting the available information in the experience while learning scalable distributed control policies.
@@ III Problem formulation
where $\mathbf{x}(t)=[(\mathbf{x}^{1}(t))^{\top},\ldots,(\mathbf{x}^{n}(t))^{\top}]\in\mathcal{X}\subseteq\mathbb{R}^{n\times n_{x}}$ is the joint state of the robot team at time $t\geq 0$, with $\mathbf{x}^{i}(t)$ the state of robot $i$ at time $t$. On the other hand, $\mathbf{u}=[(\mathbf{u}^{1}(t))^{\top},\ldots,(\mathbf{u}^{n}(t))^{\top}]\in\mathcal{U}\subseteq\mathbb{R}^{n\times n_{u}}$ is the joint input, with $\mathbf{u}^{i}(t)$ the input of robot $i$ at time $t$. The term ${\bm{\omega}}(t)$ is white noise modeling the uncertainty in the robot sensors and actuators and $\Xi=\mathbf{L}\mathbf{L}^{\top}$ is the noise diffusion matrix. Let $\{t_{\tau}\}_{\tau=0}^{\infty}$ be a sequence of discrete time instants such that $t_{\tau+1}-t_{\tau}=T>0$ and assume zero-order hold inputs such that $\mathbf{u}(t)=\mathbf{u}(t_{\tau})$, $\forall t\in[t_{\tau},t_{\tau+1})$. Then, from Eq. (1), we can obtain an Euler discretization of the dynamics with discrete-time state $\mathbf{s}_{\tau}=\mathbf{x}(t_{\tau})$, action $\mathbf{a}_{\tau}=\mathbf{u}(t_{\tau})$, and dynamics:
$\mathbf{s}_{\tau+1}=\mathbf{s}_{\tau}+T\mathbf{f}(\mathbf{s}_{\tau},\mathbf{a}_{\tau})+\mathbf{n}_{\tau},$ (2)
@@
The robots interact in a distributed manner, described by a time-varying undirected graph $\mathcal{G}_{\tau}=(\mathcal{V},\mathcal{E}_{\tau})$, where $\mathcal{E}_{\tau}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. An edge $(i,j)\in\mathcal{E}_{\tau}$ exists when robots $i$ and $j$ interact at time $\tau$. Let $\mathbf{A}_{\tau}\in\{0,1\}^{n\times n}$ be the adjacency matrix associated to $\mathcal{G}_{\tau}$, such that $[\mathbf{A}_{\tau}]_{ij}=1$ if and only if $(i,j)\in\mathcal{E}_{\tau}$, and 00 otherwise. The set of $k$-hop neighbors of robot $i$ at time $\tau$ is $\mathcal{N}^{i,k}_{\tau}=\{j\in\mathcal{V}\mid[\mathbf{A}^{k}_{\tau}]_{ij}\neq 0\}$. We remark that $\mathcal{N}^{i,k}_{\tau}$ includes robot $i$.
@@ IV-B Self-attention parameterization to enforce communication patterns
$\frac{\partial H_{{\bm{\theta}}}(\mathbf{x})}{\partial\mathbf{x}^{i}}=\sum_{j\in{\cal N}^{i,k}}\frac{\partial H^{j}_{{\bm{\theta}}}(\mathbf{x}^{i}_{{\cal N}^{j,k}})}{\partial\mathbf{x}^{i}}$ (13)
by gathering $\partial H^{j}_{{\bm{\theta}}}(\mathbf{x}^{i}_{{\cal N}^{j,k}})/\partial\mathbf{x}^{i}$ from its $k$-hop neighbors. Then, the control policy ${\bm{\mu}}_{{\bm{\theta}}}$ of robot $i$ becomes:
@@
Figure 2: Physics-informed neural network. Robot $i$ receives information from the state of its $k$-hop neighbors. Then, each self-attention-based module process the data to obtain the port-Hamiltonian terms of the controller. Finally, robot $i$ uses the learned IDA-PBC policy to compute the control input.
@@
To learn $[\mathbf{J}_{{\bm{\theta}}}(\mathbf{x})]_{ij}$, we follow the same steps in (15), with parameters $\mathbf{A}_{w}^{\mathbf{J}},\mathbf{B}_{w}^{\mathbf{J}},\mathbf{C}_{w}^{\mathbf{J}},\mathbf{D}_{w}^{\mathbf{J}}$ instead of $\mathbf{A}_{w}^{\mathbf{R}},\mathbf{B}_{w}^{\mathbf{R}},\mathbf{C}_{w}^{\mathbf{R}},\mathbf{D}_{w}^{\mathbf{R}}$, to obtain encodings $\mathbf{Z}_{ij}^{\mathbf{J}}$ instead of $\mathbf{Z}_{ij}^{\mathbf{R}}$. Due to the reciprocal communication between robots $i$ and $j$, we enforce the skew-symmetry of $\mathbf{J}_{\bm{\theta}}(\mathbf{x})$ by:
@@
Since $[\mathbf{J}_{{\bm{\theta}}}(\mathbf{x})]_{ij}=\mathbf{0}$, the interconnection matrix is such that $\mathbf{J}_{\bm{\theta}}(\mathbf{x})+\mathbf{J}_{\bm{\theta}}^{\top}(\mathbf{x})=\bf 0$ and, thus, is skew-symmetric by design. Again, each element $[\mathbf{J}_{{\bm{\theta}}}(\mathbf{x})]_{ij}$ depends only on the information from robot $i$ and its neighbor $j$, so the computation is distributed.
@@
The encodings $\mathbf{Z}^{\mathbf{M}}_{i}$ and $\mathbf{Z}^{U}_{i}$ are calculated using the same steps in (15), with parameters $\mathbf{A}_{w}^{\mathbf{M}}$, $\mathbf{B}_{w}^{\mathbf{M}}$, $\mathbf{C}_{w}^{\mathbf{M}}$, $\mathbf{D}_{w}^{\mathbf{M}}$ and $\mathbf{A}_{w}^{{U}}$, $\mathbf{B}_{w}^{{U}}$, $\mathbf{C}_{w}^{{U}}$, $\mathbf{D}_{w}^{{U}}$, respectively. Once we have $H^{i}_{{\bm{\theta}}}(\mathbf{x}^{i}_{{\cal N}^{i,k}})$, we obtain $\partial H^{j}_{{\bm{\theta}}}(\mathbf{x}^{i}_{{\cal N}^{j,k}})/\partial\mathbf{x}^{i}$ from all the neighboring robots and compute Eq. (13). Therefore, since all the operations are only dependent on the available information in robot $i$ and its neighbor $j$, the computations involving the Hamiltonian function are distributed by design.
@@ V Physics-informed multi-robot soft actor-critic
Figure 3: Overview of our physics-informed multi-agent reinforcement learning approach. The soft actor-critic method consists of: (a) an actor, the IDA-PBC self-attention-based neural network policy, (b) an interactive environment with, e.g., goals to reach and adversaries to avoid, and that keeps track of the correlations among robots encoded in the communication graph, (c) a global reward function that describes the task and steers the multi-agent reinforcement learning training without any particular factorization, and (d) a centralized critic, modeled as a neural network that learns the state-action value function relating the agent and the environment according to the desired task, which allows to learn the same policy for all the robots simultaneously with the same policy and value updates.
@@ V-A The actor
$\mathbf{a}_{\tau}^{i}\sim{\bm{\pi}}_{\bm{\theta}}(\mathbf{a}_{\tau}^{i}|\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})=\tanh({\bm{\mu}}_{\bm{\theta}}(\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})+{\bm{\sigma}}_{\bm{\psi}}(\mathbf{a}_{\tau}^{i},\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})\xi),$ (19)
where ${\bm{\mu}}_{\bm{\theta}}(\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})$ is the IDA-PBC policy parameterized by self-attention neural networks under the restrictions imposed by the networked structure of the robot team. On the other hand, ${\bm{\sigma}}_{\bm{\psi}}(\mathbf{a}_{\tau}^{i},\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})$ is a vector of standard deviations given by a neural network that approximates the variance of the squashed Gaussian distribution with parameters ${\bm{\psi}}$. Besides, $\xi\sim{\cal N}(\mathbf{0},\mathbf{I})$, so the control policy is a Gaussian distribution with mean ${\bm{\mu}}_{\bm{\theta}}(\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})$ and diagonal covariance matrix whose diagonal is equal to $({\bm{\sigma}}_{\bm{\psi}}(\mathbf{a}_{\tau}^{i},\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}}))^{2}$. The control is constrained to $\mathbf{a}_{\tau}^{i}\in[-1,1]^{n_{u}}$ by means of a $\tanh$ function [75, 54, 73], leading to a squashed Gaussian policy. In practice, the control input can be constrained to $\mathbf{a}_{\tau}^{i}\in[a_{\min},a_{\max}]^{n_{u}}$ with $-\infty<a_{\min}\leq a_{\max}<\infty$ by scaling the output of the $\tanh$ function. It is also important to remark that ${\bm{\sigma}}_{\bm{\psi}}(\mathbf{a}_{\tau}^{i},\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})$ only depends on the available information at each robot, such that the desired distributed factorization of the control policy is preserved.
During training, the control inputs are sampled from the Gaussian distribution, as stated in Eq. (3). After training, the mean control input is chosen, i.e., $\mathbf{a}^{i}_{\tau}=\tanh({\bm{\mu}}_{\bm{\theta}}(\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})).$ In the case that an additional layer of robust, adaptive or active control is desired, then ${\bm{\sigma}}_{\bm{\psi}}(\mathbf{a}_{\tau}^{i},\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})$ can be used as a proxy of the uncertainty in the control policy, since it only depends on the neighboring information at robot $i$. The design of ${\bm{\sigma}}_{\bm{\psi}}(\mathbf{a}_{\tau}^{i},\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})$ is free to choose, but in this work we opt for the same architecture in Eq. (15).

### v3

@@ Abstract
The networked nature of multi-robot systems presents challenges in the context of multi-agent reinforcement learning. Centralized control policies do not scale with increasing numbers of robots, whereas independent control policies do not exploit the information provided by other robots, exhibiting poor performance in cooperative-competitive tasks. In this work we propose a physics-informed reinforcement learning approach able to learn distributed multi-robot control policies that are both scalable and make use of all the available information to each robot. Our approach has three key characteristics. First, it imposes a port-Hamiltonian structure on the policy representation, respecting energy conservation properties of physical robot systems and the networked nature of robot team interactions. Second, it uses self-attention to ensure a sparse policy representation able to handle time-varying information at each robot from the interaction graph. Third, we present a soft actor-critic reinforcement learning algorithm parameterized by our self-attention port-Hamiltonian control policy, which accounts for the correlation among robots during training while overcoming the need of value function factorization. Extensive simulations in different multi-robot scenarios demonstrate the success of the proposed approach, surpassing previous multi-robot reinforcement learning solutions in scalability, while achieving similar or superior performance (with averaged cumulative reward up to ×$\times 2$ greater than the state-of-the-art with robot teams ×$\times 6$ larger than the number of robots at training time). We also validate our approach on multiple real robots in the Georgia Tech Robotarium under imperfect communication, demonstrating zero-shot sim-to-real transfer and scalability across number of robots.
@@ I Introduction
(a) navigation (b) sampling
Refer to caption Refer to caption
(c) transport (d) grassland
Refer to caption Refer to caption
(e) 6x1-Half Cheetah
(f) multi-robot Robotarium [1] navigation
Refer to caption Refer to caption
Figure 1: Examples of scenarios addressed by our physics-informed multi-agent reinforcement learning approach. The scenarios cover a wide variety of cooperative/competitive behaviors and levels of coordination complexity: (a) robots learn to navigate to landmarks (black dots) while avoiding collisions with other robots (colored trajectories from pink at $t=0$ to blue at $t=t_{f}$); (b) robots cooperate to do active LiDAR (star pattern) sampling (orange areas refer to regions already sampled) in an unknown environment with regions of low reward (magenta areas) and high reward (green areas); (c) robots (blue dots) cooperate to transport a square box (red) with unknown mass and that follows Newtonian dynamics towards a desired region (green dot, box turns green when it reaches the goal); (d) the robots (blue) collect food (green) while avoiding collisions with attackers (red), since collision with attackers lead to deactivation of the robots; (e) joints learn how to coordinate to make the Half-Cheetah robot walk, where each joint is an independent agent that can only interact with the adjacent agents (image from [2]); (f) a team of robots learns to navigate and avoid collisions in a real-world setting characterized by a tight space and high density of robots.
@@
Extensive simulations and real experiments in multi-robot scenarios (Sec. VI) demonstrate that the combination of multi-agent reinforcement learning techniques and a physics-informed description of the system achieves scalable control policies with similar or superior performance as the state-of-the-art (Sec. VII).
This paper is an extension of our prior conference paper [19]. The main difference of this work compared to [19] is that it addresses a multi-agent reinforcement learning problem (where only reward samples are provided), whereas [19] solves a learning-from-demonstrations problem (where data from an expert control policy is available). This leads to a new multi-agent reinforcement learning problem formulation and a substantially different approach beyond the policy parameterization, since our new approach involves modifications on the parameterization of the actor and the replay buffer of the actor-critic, besides some particularities in the reward function design and the critic. The neural networks that model the control policies are different as well, since the multi-agent reinforcement learning scenarios are continuous-time while those in [19] are discrete-time. Finally, all experiments in this paper are new, including an extensive scalability analysis in a real multi-robot platform involving up to $16$ robots in a tight space.
@@ II-B Multi-agent reinforcement learning for robotic problems
To solve the non-stationarity issue, a common approach is to factorize the value and/or policy functions. The most common factorization, the so-called centralized-training decentralized-execution, departs from centralized-training reinforcement learning approaches [56, 57, 58] and considers that the agents act independently. To provide some global information to the agents, agents learn an approximation of the other agents’ control policies [59, 60, 61]. However, this proxy is only used during training and, since the resulting control policies only use the individual observations, the performance deteriorates the number of robots change. This aspect is partially overcome by considering heterogeneous agents [62] but scalability is deteriorated again in large teams because the agents do not learn how to exploit neighboring information effectively. Another option is to provide each agent with a global estimator that predicts the trajectories of the whole multi-robot team [63] or to implicitly learn to coordinate with other agents through the value function [55]. In mean field approaches [64], the multi-agent game is reduced to an interaction between an agent and the average of the other agents, which is not practical. To overcome these issues, we propose a new formulation of the policy function that is distributed by design and captures the underlying interaction graph that describes the multi-robot system.
Other works focus on more complex factorizations that account for local interactions among neighboring agents in the value function. For instance, the value function can be approximated as dependent on the neighbors only [65]. Either using analytical [35, 16, 66] or learned [67, 68, 69] factorizations, the control policies are still restricted to the same team size used during training because the policies do not explicitly consider communication. Closely related to ours, the authors of [70] propose an end-to-end neural network based on attention to achieve scalability of the control policy in quadrotor swarm navigation problems. Similarly, by using a port-Hamiltonian description of the system, our approach avoids factorizing the reward or the value functions, thus fully exploiting the available information in the experience while learning scalable distributed control policies.
@@ III Problem formulation
where $\mathbf{x}(t)=[(\mathbf{x}^{1}(t))^{\top},\ldots,(\mathbf{x}^{n}(t))^{\top}]\in\mathcal{X}\subseteq\mathbb{R}^{n\times n_{x}}$ is the joint state of the robot team at time $t\geq 0$, with $\mathbf{x}^{i}(t)$ the state of robot $i$ at time $t$. On the other hand, $\mathbf{u}=[(\mathbf{u}^{1}(t))^{\top},\ldots,(\mathbf{u}^{n}(t))^{\top}]\in\mathcal{U}\subseteq\mathbb{R}^{n\times n_{u}}$ is the joint input, with $\mathbf{u}^{i}(t)$ the input of robot $i$ at time $t$. The term ${\bm{\omega}}(t)$ is white noise modeling the uncertainty in the robot sensors and actuators and $\Xi=\mathbf{L}\mathbf{L}^{\top}$ is the noise diffusion matrix. Let $\{t_{\tau}\}_{\tau=0}^{\infty}$ be a sequence of discrete time instants such that $t_{\tau+1}-t_{\tau}=T_{\tau}>0$ and assume zero-order hold inputs such that $\mathbf{u}(t)=\mathbf{u}(t_{\tau})$, $\forall t\in[t_{\tau},t_{\tau+1})$. Then, from Eq. (1), we can obtain an Euler discretization of the dynamics with discrete-time state $\mathbf{s}_{\tau}=\mathbf{x}(t_{\tau})$, action $\mathbf{a}_{\tau}=\mathbf{u}(t_{\tau})$, and dynamics:
$\mathbf{s}_{\tau+1}=\mathbf{s}_{\tau}+T_{\tau}\mathbf{f}(\mathbf{s}_{\tau},\mathbf{a}_{\tau})+\mathbf{n}_{\tau},$ (2)
@@
The robots interact in a distributed manner, described by a time-varying undirected graph $\mathcal{G}_{\tau}=(\mathcal{V},\mathcal{E}_{\tau})$, where $\mathcal{E}_{\tau}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. An edge $(i,j)\in\mathcal{E}_{\tau}$ exists when robots $i$ and $j$ interact at time $\tau$. Let $\mathbf{A}_{\tau}\in\{0,1\}^{n\times n}$ be the adjacency matrix associated to $\mathcal{G}_{\tau}$, such that $[\mathbf{A}_{\tau}]_{ij}=1$ if and only if $(i,j)\in\mathcal{E}_{\tau}$, and 00 otherwise. The set of $k$-hop neighbors of robot $i$ at time $\tau$ is $\mathcal{N}^{i,k}_{\tau}=\{j\in\mathcal{V}\mid[\mathbf{A}^{k}_{\tau}]_{ij}\neq 0\}$, where $\mathbf{A}^{k}_{\tau}$ is the $k$-th power of matrix $\mathbf{A}_{\tau}$. We remark that $\mathcal{N}^{i,k}_{\tau}$ includes robot $i$.
@@ IV-B Self-attention parameterization to enforce communication patterns
$\frac{\partial H_{{\bm{\theta}}}(\mathbf{x})}{\partial\mathbf{x}^{i}}=\sum_{j\in{\cal N}^{i,k}}\frac{\partial H^{j}_{{\bm{\theta}}}(\mathbf{x}^{j}_{{\cal N}^{j,k}})}{\partial\mathbf{x}^{i}}$ (13)
by gathering $\partial H^{j}_{{\bm{\theta}}}(\mathbf{x}^{j}_{{\cal N}^{j,k}})/\partial\mathbf{x}^{i}$ from its $k$-hop neighbors. Then, the control policy ${\bm{\mu}}_{{\bm{\theta}}}$ of robot $i$ becomes:
@@
Figure 2: Physics-informed policy parameterization. At each instant, robot $i$ receives state information from its neighbors, typically associated to a perception or communication range. Then, three self-attention-based modules, each one associated to a component of the desired closed-loop port-Hamiltonian dynamics ($[\mathbf{R}_{\theta}(\mathbf{x})]_{ij}$, $[\mathbf{J}_{\theta}(\mathbf{x})]_{ij}$, $H_{\theta}^{i}(\mathbf{x}$)), use the information from the neighbors to compute the parameters of an interconnection and damping assignment passivity-based control (IDA-PBC) policy, which is then used to execute the desired control action.
@@
To learn $[\mathbf{J}_{{\bm{\theta}}}(\mathbf{x})]_{ij}$, we follow the same steps in (LABEL:eq:R_layers), with parameters $\mathbf{A}_{w}^{\mathbf{J}},\mathbf{B}_{w}^{\mathbf{J}},\mathbf{C}_{w}^{\mathbf{J}},\mathbf{D}_{w}^{\mathbf{J}}$ instead of $\mathbf{A}_{w}^{\mathbf{R}},\mathbf{B}_{w}^{\mathbf{R}},\mathbf{C}_{w}^{\mathbf{R}},\mathbf{D}_{w}^{\mathbf{R}}$, to obtain encodings $\mathbf{Z}_{ij}^{\mathbf{J}}$ instead of $\mathbf{Z}_{ij}^{\mathbf{R}}$. Due to the reciprocal communication between robots $i$ and $j$, we enforce the skew-symmetry of $\mathbf{J}_{\bm{\theta}}(\mathbf{x})$ by:
@@
Since $[\mathbf{J}_{{\bm{\theta}}}(\mathbf{x})]_{ii}=\mathbf{0}$, the interconnection matrix is such that $\mathbf{J}_{\bm{\theta}}(\mathbf{x})+\mathbf{J}_{\bm{\theta}}^{\top}(\mathbf{x})=\bf 0$ and, thus, is skew-symmetric by design. Again, each element $[\mathbf{J}_{{\bm{\theta}}}(\mathbf{x})]_{ij}$ depends only on the information from robot $i$ and its neighbor $j$, so the computation is distributed.
@@
The encodings $\mathbf{Z}^{\mathbf{M}}_{i}$ and $\mathbf{Z}^{U}_{i}$ are calculated using the same steps in (LABEL:eq:R_layers), with parameters $\mathbf{A}_{w}^{\mathbf{M}}$, $\mathbf{B}_{w}^{\mathbf{M}}$, $\mathbf{C}_{w}^{\mathbf{M}}$, $\mathbf{D}_{w}^{\mathbf{M}}$ and $\mathbf{A}_{w}^{{U}}$, $\mathbf{B}_{w}^{{U}}$, $\mathbf{C}_{w}^{{U}}$, $\mathbf{D}_{w}^{{U}}$, respectively. Once we have $H^{i}_{{\bm{\theta}}}(\mathbf{x}^{i}_{{\cal N}^{i,k}})$, we obtain $\partial H^{j}_{{\bm{\theta}}}(\mathbf{x}^{j}_{{\cal N}^{j,k}})/\partial\mathbf{x}^{i}$ from all the neighboring robots and compute Eq. (13). Therefore, since all the operations are only dependent on the available information in robot $i$ and its neighbor $j$, the computations involving the Hamiltonian function are distributed by design.
@@ V Physics-informed multi-robot soft actor-critic
Figure 3: Overview of our physics-informed soft actor-critic multi-agent reinforcement learning approach. The main differences with respect of other actor-critic methods are the following: (a) the actor is the multi-robot network modeled as a port-Hamiltonian system with a self-attention-based IDA-PBC policy; (b) the replay buffer not only stores actions, states and rewards, but also the graph structure of the multi-robot network to enforce the desired distributed structure; (c) the reward is global because the actor is the whole multi-robot team and the physics-informed parameterization conditions the policy on the graph structure of the team; (d) the output of the critic is shared across robots and the policy parameters are the same for all robots, so for the same critic gradient step, $n$ policy gradient steps are taken.
@@ V-A The actor
$\mathbf{a}_{\tau}^{i}\sim{\bm{\pi}}_{\bm{\theta}}(\mathbf{a}_{\tau}^{i}|\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})=\tanh({\bm{\mu}}_{\bm{\theta}}(\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})+{\bm{\sigma}}_{\varrho}(\mathbf{a}_{\tau}^{i},\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})\xi),$ (19)
where ${\bm{\mu}}_{\bm{\theta}}(\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})$ is the IDA-PBC policy parameterized by self-attention neural networks under the restrictions imposed by the networked structure of the robot team.
On the other hand, ${\bm{\sigma}}_{\varrho}(\mathbf{a}_{\tau}^{i},\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})$ is a vector of standard deviations given by a neural network that approximates the variance of the squashed Gaussian distribution with parameters $\varrho$. Besides, $\xi\sim{\cal N}(\mathbf{0},\mathbf{I})$, so the control policy is a Gaussian distribution with mean ${\bm{\mu}}_{\bm{\theta}}(\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})$ and diagonal covariance matrix whose diagonal is equal to $({\bm{\sigma}}_{\varrho}(\mathbf{a}_{\tau}^{i},\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}}))^{2}$. The control is constrained to $\mathbf{a}_{\tau}^{i}\in[-1,1]^{n_{u}}$ by means of a $\tanh$ function [82, 56, 80], leading to a squashed Gaussian policy. In practice, the control input can be constrained to $\mathbf{a}_{\tau}^{i}\in[a_{\min},a_{\max}]^{n_{u}}$ with $-\infty<a_{\min}\leq a_{\max}<\infty$ by scaling the output of the $\tanh$ function. It is also important to remark that ${\bm{\sigma}}_{\varrho}(\mathbf{a}_{\tau}^{i},\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})$ only depends on the available information at each robot, such that the desired distributed factorization of the control policy is preserved.
During training, the control inputs are sampled from the Gaussian distribution, as stated in Eq. (3). After training, the mean control input is chosen, i.e., $\mathbf{a}^{i}_{\tau}=\tanh({\bm{\mu}}_{\bm{\theta}}(\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})).$ In the case that an additional layer of robust, adaptive or active control is desired, then ${\bm{\sigma}}_{\varrho}(\mathbf{a}_{\tau}^{i},\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})$ can be used as a proxy of the uncertainty in the control policy, since it only depends on the neighboring information at robot $i$. The design of ${\bm{\sigma}}_{\varrho}(\mathbf{a}_{\tau}^{i},\mathbf{s}^{i}_{{\cal N}_{\tau}^{i,k}})$ is free to choose, but in this work we opt for the same architecture in Eq. (LABEL:eq:R_layers).